# Transfer Learning with Foundational Models for Time Series Forecasting using Low-Rank Adaptations

### v1

@@ 4.1 Experimental seutp
In the experimental study, we also observed the training time and the missing rate [46]. Given that LLMs are not inherently designed for TSF tasks, we must account for their potential to generate incorrect outputs, known as hallucinations or anomalies. In our work we classify an output as an anomaly if it fails to forecast a horizon shorter than the expected one for each dataset. If the model outputs a longer horizon, we will use only the first $h$ values and disregard the rest. The missing rate (MR) is defined in equation 7 as the difference between the number of set instances $n_{test}$ and the correctly outputted ones $n_{decoded}$ divided by the total size of the test set multiplied by 100.
$MissingRate\%=\frac{n_{test}-n_{decoded}}{n_{test}}\cdot 100$ (7)
4.2 Experimental study
In this study, it is tested how LLIAM is able to emit forecasts whose quality is competent in comparison with the average case presented in conventional DL models, such as RNNs and TCNs.
The methodology employed is delineated as follows. Initially, an instance of LLIAM is trained with the specified hyperparameters, as presented in Table 2. It is crucial to highlight that all seven datasets are consolidated into a unified dataset, which is employed for fine-tuning the proposed approach. In contrast, conventional DL methods, such as RNNs and TCNs, must be trained on a dataset-by-dataset basis, with all combinations of hyperparameters listed in Table 4. The metrics presented in the tables related to these algorithms are the averages of all configurations. The objective of this study is to demonstrate that LLIAM is capable of achieving performance levels that are comparable to or superior to those of these methods when a baseline hyperparameter configuration is employed. The combination of all datasets is intended to demonstrate that the comprehensive knowledge inherent to FM models, such as LLMs, enables the development of general models, reducing the necessity for time-consuming hyperparameter optimization for each dataset.
Prior to the construction of the sliding windows, a preprocessing pipeline was applied to the seven datasets. First, the heuristic methodology described in [57] is employed for the purposes of anomaly detection and mitigation. For the conventional DL methods, a min-max normalization is also applied over the full dataset to ensure the convergence of these models. After constructing the windows for each series of the dataset, the training and test sets are constructed using a leave-one-out strategy, with the last generated window of each series being used for testing. Consequently, the size of the test set is equal to the number of series in the dataset.
The results confirm our hypothesis regarding the generalization capabilities of fine-tuned LLMs for the univariate TSF task. As can be seen in Table 5, the differences between LLIAM and the other models are significant, with lower RMSE at both temperatures. For example, in the Electricity dataset, LLIAM T=10 and T=20 achieve RMSE values of 45120.043 and 42338.992 respectively, significantly outperforming the second-best model, LSTM, which achieves an RMSE of 287414.428. For the M1 monthly and quarterly data sets, LLIAM T=10 also leads with RMSE values of 2464.424 and 1913.682, outperforming the conventional DL models. The same occurs with the rest of the datasets, with our proposal achieving the lowest of the results 4 times with T=10 and three times with T=20. While the RMSE values in some datasets might seem high, they reflect the larger scale and range of the data in those specific cases. Also, it is important to note that RMSE is sensitive to anomalies. The consistently lower RMSE values of LLIAM across different datasets indicate its robustness and superior performance in the TSF task.
Unlike RMSE, SMAPE takes into account the scale of the data and determines whether a model can make predictions within the desired range. The results reported in Table 6 inform us that our propoasl. LLIAM achieves the best performance against conventional algorithms for each dataset and globally (as shown in the Avg. of Avg. column). As can be seen, compared to the best non-foundational DL model, we are able to obtain a SMAPE of 0.997 with LLIAM T=20, respecting a SMAPE of 0.232 given by the TCN D=0.1. In M1 Monthly, this performance improvement is pronounced, going from 0.436 of the TCN D=0 to 0.163 with LLIAM T=10. The M1 Quarterly results show the same tendency, obtaining a reduction of the error from 0.481 to 0.171 with LLIAM T=10. However, on the M3 and NN5 datasets the error commitment is similar between LLIAM and the other methods, achieving a small error reduction on the M3 Monthly and NN5 datasets and matching it on the M3 Quarterly with LSTM and TCN D=0 models. Overall, the performance of LLIAM T=10 and T=20 is similar, with an average SMAPE of 0.147 and 0.150 respectively, significantly outperforming all other methods.
One aspect to consider is the ability of our proposal to produce coherent and expected responses. It must be able to predict at any given time without resorting to making up hallucinations. The missing rate for each dataset and LLIAM configuration is reported in Table 7. Hallucinations were observed only on NN5 Daily for LLIAM T=10 and also on M1 Monthly for LLIAM T=20. The most frequent anomaly is due to the model stopping early, failing to reach the expected forecast horizon. Early studies showed higher missing rates, but increasing the forecast horizon by one step ($h+1$) reduced this issue. Overall, the hallucinations are not significant with respect to the results obtained.
Finally, it is necessary to consider the training time of LLIAM compared to the other models, as detailed in Table 8. Since temperature is only an inference parameter, our proposal only required one training phase. Additionally, kindly note that this only instance was trained over all the datasets with its default hyperparameters unlike LSTM, GRU and TCNs. The total time for these three methods required an exhaustive set of experiments for each configuration and dataset, resulting in a total time that exceeds the time required to train LLIAM (about 3 days compared to 18 days in the best case). Its performance suggests that fine-tuning pre-trained FMs may lead to the creation of more sustainable proposals compared to training a specific model from scratch.
@@ 4.2 Experimental study
LLIAM T=10 45120,043 2464,424 1913,682 807,877 629,930 5,918 18,342
LLIAM T=20 42338,992 2348,506 2655,757 829,951 631,046 5,906 19,085
GRU 436904,097 27405,683 29961,260 1335,165 1028,065 6,887 24,109
LSTM 287414,428 24225,724 21713,865 1265,023 964,428 6,677 22,682
TCN D=0 670686,818 34761,041 13997,682 1326,214 973,804 6,633 22,704
TCN D=0.1 663362,613 34764,894 15835,914 1347,495 1003,729 6,712 22,721
TCN D=0.25 683029,861 32666,035 17290,540 1368,059 1001,894 6,863 23,111
Table 5: RMSE metric per model and dataset. LSTM, GRU and TCN RMSE are averages for all the configurations tested, meanwhile LLIAM is the average RMSE obtained after aggregating the test instances.
@@
of Avg.
LLIAM T=10 0,099 0,163 0,171 0,150 0,102 0,232 0,114 0,147
LLIAM T=20 0,097 0,166 0,175 0,153 0,103 0,237 0,119 0,150
GRU 0,241 0,604 0,616 0,165 0,113 0,259 0,135 0,305
LSTM 0,239 0,624 0,579 0,153 0,102 0,247 0,127 0,296
TCN D=0 0,242 0,436 0,481 0,160 0,102 0,246 0,129 0,257
TCN D=0.1 0,232 0,471 0,485 0,166 0,111 0,249 0,129 0,263
TCN D=0.25 0,247 0,478 0,533 0,168 0,108 0,257 0,132 0,275
Table 6: SMAPE metric per model and dataset. LSTM, GRU and TCN SMAPE are averages for all the configurations tested, meanwhile LLIAM is the average SMAPE obtained after aggregating the test instances.
@@
LLIAM T=10 0,000 0,000 0,000 0,000 0,000 3,604 0,000 0,5148
LLIAM T=20 0,000 0,486 0,000 0,000 0,000 6,306 0,000 0,9704
Table 7: Percentage of test instances unable to decode into the expected forecast horizon. GRU, LSTM and TCN models have been ommited because their missing rate is always 0%.
Total training time in seconds (Lower is better)
LLIAM LSTM GRU
TCN
D=0
TCN
D=0.1
TCN
D=0.25
266563,780 s 1612037,888 s 1648122,031 s 1819626,881 s 1899311,350 s 1883758,756 s
3d 2h 2m 43s 18d 15h 47m 18s 19d 1h 48m 42 s 21d 1h 27m 7s 21d 23h 35m 11s 21d 19h 15m 59s
Table 8: Total training time in seconds. GRU, LSTM and TCNs training times are the sum of all the experiments made for all the configurations, meanwhile only one instance of LLIAM has been trained over all the datasets. A conversion from seconds to days (d), hours (h), minutes (m) and seconds (s) is also supplied.
4.3 Zero-shot study
In the zero-shot study we show that our adaptation of the model achieves a good degree of generalization not only for known domains, but also for unknown ones, proving that LLMs can be zero-shot forecasters as stated in [5]. The remaining 3 datasets (San Francisco traffic, ETTh1 and ETTh2) are used in this study. It is important to note that the domain of the San Francisco traffic dataset does not resemble any of the domains from the dataset used to build LLIAM. Due to the large length of the ETTh1 and ETTh2 datasets, we only used the last 10% of the generated windows to test the models. The LLaMA-1 7B base architecture is used as the baseline model, with the same temperature values used in LLIAM.
Table 9 highlights the average RMSE and SMAPE obtained for each model configuration. In ETTh1 and San Francisco traffic, the results obtained with LLIAM surpass the performance of the baseline model. However, the error between our proposal and the base model is very similar in the ETTh2 dataset. Attention may be drawn by the high average error obtained with LLAMA T=20 in the dataset ETTh1. This is produced by an anomalous behavior, where the model predicted a high value in a test window whose tendency is led by low ones. We assume that it occurred as the temperature of the model increased, since it does not occur at a value of T=10. In addition, it appears that fine-tuning with time series can also mitigate this behavior, as it did not occur when using LLIAM. Evaluating the results obtained according to the SMAPE metric, we can conclude that both metrics reach a consensus, as in the previous study. LLIAM T=10 presents a lower average SMAPE with a value of 0.171, followed by LLIAM T=20 with a similar one. This also occurs with LLAMA, obtaining an error of 0.190 with T=10 and 0.188 with T=20.
RMSE (Lower is better) SMAPE (Lower is better)
ETTh1 ETTh2 Traffic ETTh1 ETTh2 Traffic Average SMAPE
LLIAM T=10 2,053 8,014 1,550 0,184 0,199 0,129 0,171
LLIAM T=20 2,127 8,016 1,600 0,190 0,198 0,134 0,174
LLAMA T=10 2,396 7,768 1,827 0,213 0,198 0,158 0,190
LLAMA T=20 3822528,513 7,562 1,845 0,211 0,194 0,159 0,188
@@ 4.3 Zero-shot study
5 Final discussions
This proposal introduces a novel approach to the TSF task using FMs, specifically the LLaMA model. It combines the PEFT technique LoRA with a modified PromptCast method for inputting time series data into the model. In addition, two sets of experiments have been conducted: one regarding the effectiveness and efficiency of this approach relative to others, and another to validate and compare the use of these models for TSF on unseen data sets.
The results obtained are competent and verify the outperforming of LLIAM over the other DL non-foundational methods, being capable of giving forecast with a better or equal quality than them. The intrinsic properties of FMs, highlighting it great generalization capacity by adopting TL as one of its key qualities, and the effectiveness of the fine-tuning technique used has led to the development of our proposal achieving its positive outcomes. In addition, it is noteworthy that we trained the model only once, especially considering that the training time is significantly lower compared to the total time required to train all configurations of the LSTM, GRU, and TCN models. Therefore, it is clear to say that LLIAM is more efficient.
Throughout our research, we have discussed how temperature affects the quality of the forecasts issued. As explained in Section 4.1, this hyperparameter affects the randomness of the model output. Consequently, we have only tested with low temperature values in order to respect the original non-normalized probability distribution emitted by the model. This was done to avoid unusual predictions and to encourage more deterministic ones. Light evidence has been found verifying that lower temperature values encourage better results.
The use of PEFT techniques to fine-tune pre-trained models has been widely applied to text, image, and even audio generation with good results. The sequential nature of text and audio data inspired us to try to adapt LLMs to the TSF task. As demonstrated, this initial work proved that FMs can be adequate for performing this task. However, in the case of LLMs, it is necessary to perform a tedious preliminary work to determine a number of factors related to how the data is input into them and how to adapt them. Another interesting topic to be researched is in the minimization of the missing rate, which is required to guarantee robustness with respect to the behavior of these models. Nevertheless, the opportunities offered by these types of models are extensive and constitute an interesting line of future work regarding their adaptation to different tasks as fine-tuned or zero-shot predictors.
This positive initial study marks the beginning of future work involving more extensive testing with different models of a similar nature, where they can be adapted or built specifically for TSF-related tasks. Actually, few attempts are found related to the adaptation of these FMs to these types of tasks. In addition, the exploration and creation of diverse PEFT techniques that go beyond LoRAs is also an interesting and wide-ranging aspect of the effective and efficient use and tuning of this type of models. Overall, the arrival of the FMs opens up the possibility of numerous interesting proposals.

### v3

@@ 4.1 Experimental setup
In the experimental study, the missing rate (MR) [40] is also observed. Given that LLMs are not inherently designed for TSF tasks, it must be taken into account for their potential to generate incorrect outputs, known as hallucinations or anomalies. In our work an output with a horizon shorter than the expected one for each dataset is classified as an anomaly. If the model outputs a longer horizon, a trim of the first $h$ values is applied and the rest are disregarded. The missing rate is defined in equation 7 as the difference between the number of set instances $n_{test}$ and the correctly outputted ones $n_{decoded}$ divided by the total size of the test set multiplied by 100.
$MR=\frac{n_{test}-n_{decoded}}{n_{test}}\cdot 100$ (7)
4.2 Comparative study
In this study our proposal, LLIAM, is compared with various conventional DL algorithms, such as RNNs and TCNs, and TimeLLM, a state-of-the-art reprogramming framework to enable the forecasting of time-series with LLMs presented in Section 2.1. The present study aims to provide evidence that the general training approach employed in LLIAM surpasses the traditional dataset-specific training paradigm. In contrast to other existing methods, such as TimeLLM, which do not support this level of generalization, LLIAM is designed to capture broader patterns across all datasets.
Methodology
We have grouped the methodology employed for each type of model evaluated in this study, as not all of them undergo the same training process.
  • 1.
    Traditional DL algorithms (RNNs, TCNs) are trained on a dataset-by-dataset basis with all combinations of hyperparameters listed in Table 4, being the metrics presented in the tables related to these algorithms averages of all configurations. The purpose of doing this is to be able to obtain a value that corresponds to an average (or ”common”) case of these algorithms.
  • 2.
    TimeLLM is also trained on a dataset-by-dataset basis due to the presence of a final lineal layer that projects the codification emitted by the LLM into the size of the expected horizon [37]. The results presented correspond to the average values of all test instances and not different configurations. We have not performed any hyperparameter optimization because it is our interest to find if a baseline configuration of a LLM-based model can surpass the conventional DL algorithms. We have used the same hyperparameters as in [37] for the Weather and ILI datasets when training TimeLLM. Otherwise, the same hyperparameters as for the M4 competition are used. The number of epochs has been set to 10, given by an heuristic. For our study, the validation set was removed and included in the training set.
  • 3.
    LLIAM is trained only one time, combining all train partitions into a whole training dataset. As elements that depend on the length of the series to emit the forecast (i.e. a lineal layer with $H$ units) are not included, a more general model than the DL ones and TimeLLM is builded. TimeLLM has not been trained this way as it is not supported by their architecture, although it is a LLM-based model. Following the premise mentioned earlier, hyperparameter optimization is not performed over LLIAM. We have used a baseline configuration specified in Section 3.
All of the series have been preprocessed for the detection and mitigation of anomalous values using the heuristic described in [56]. For the RNNs and TCNs, a Min-Max normalization is applied to incentive the convergence into a local-optima of these methods. For TimeLLM a standardization of the data is done, as seen in its GitHub repository [^2] https://github.com/KimMeen/Time-LLM, and a brief description of each dataset has also been incorporated because it is required by its prompting scheme. Both normalization and standarization is reversed when calculating the corresponding metrics. Standardization or normalization techniques on the values of the series are not necessary for LLIAM nor dataset-related information is included.
The construction of the training and test partitions is straightforward. For datasets with multiple series the sliding window methodology described in Subsection 4.1 is applied. Then, a leave-one-out strategy is employed over Electricity, M1 Monthly, M1 Quarterly, M3 Monthly, M3 Quarterly, NN5 Daily and NN5 Weekly. For the remaining two datasets (Weather and ILI) the number of instant contained in the test partition from [37] is extracted and after that, the sliding window strategy is applied, obtaining the test and train windows. Weather test set has been shortened to the last 10% windows due to its size.
Results
The results obtained confirmed that LLM-based techniques, LLIAM and TimeLLM, outperforms RNNs and TCNs in every dataset. As seen in Tables 5 and 6, the difference between these techniques average case and LLM-based ones are significant for both metrics. This suggests that LLMs may be capable of understanding temporal patterns in time-series data. It is important to note that LLIAM does not require the inclusion of any additional domain-specific or dataset-related information, such as univariate statistics, as is necessary for TimeLLM. Furthermore, it does not necessitate the standardization or normalization of the data as is required by all the other models, which serves to reinforce the veracity of this hypothesis.
When comparing both LLM-based methods using RSME (Table 5) it is easily noticed that both achieve similar performance except on the Electricity dataset. The outstanding results obtained by LLIAM and TimeLLM inform us that the generalization enabled by Transfer Learning and the similitude between textual and time-series data allow LLM models to perform the TSF task, activating the time-series analysis knowledge presented in these type of FMs improved by fine-tuning. It is important to note that RMSE, while commonly used, does not allow for direct comparisons across multiple datasets and is not appropriate for statistical testing unlike SMAPE. Furthermore, it is noteworthy that LLIAM possesses the capacity to execute generic training with all datasets combined, without the necessity for additional information or pre-processing steps, in contrast to the limitations exhibited by TimeLLM.
Table 6 presents the results calculated using SMAPE. The results obtained form LLIAM and TimeLLM are comparable across the majority of the data sets. Additionally, in both M1 Monthly and M1 Quarterly, LLIAM scores consistently achive results that are twice as high (in precision) as TimeLLM. In the column ’Avg. of Avgs.’ is presented how that LLIAM with T=10 achieves the best result followed by LLIAM T=20 and then TimeLLM. This achievement reinforces the hypothesis that it is not necessary to implement complex adaptation mechanisms on LLMs to achieve good results.
To consolidate this analysis, two non-parametric statistical tests using the SMAPE metric have been performed. First, a Friedman test is done, demonstrating that there are significant differences between the results obtained from the different methods evaluated. Then, a paired Wilcoxon test is performed to detect statistical differences between each pair of methods. A critical differences diagram has been made to illustrate better the results obtained after performing these tests. Figure 3 illustrates the post-hoc analysis of the Wilcoxon test performed. As expected, no statistical difference has been found between the different LLM-based proposals although a clear gap is set between the conventional DL ones. LLIAM T=10 obtains the highest rank, followed by TimeLLM and LLIAM T=20 which indicates that the adaptation techniques used are capable of achieving competent results.
@@ Results
ILI Weather
LLIAM T=10 55293.642 2715.727 2968.888 803.134 634.164 6.279 18.468 226836.515 9.204
LLIAM T=20 53743.873 2347.011 2644.296 818.799 639.219 6.296 19.399 217596.959 8.580
TimeLLM 36781.131 2920.872 2889.753 801.426 602.462 5.677 18.789 171705.697 10.148
GRU 436904.097 27405.683 29961.260 1335.165 1028.065 6.887 24.109 493659.388 18.204
LSTM 287614.428 24225.724 21713.865 1265.023 964.428 6.677 22.682 492403.464 18.860
TCN D=0 670686.818 34761.041 13397.682 1326.214 973.804 6.633 22.704 304509.508 21.276
TCN D=0.1 663362.613 34764.894 15835.914 1347.495 1003.729 6.712 22.721 306990.855 21.128
TCN D=0.25 683029.861 32666.035 17290.540 1368.059 1001.894 6.863 23.111 - -
Table 5: RMSE metric per model and dataset. LSTM, GRU and TCN RMSE are averages for all the configurations tested, meanwhile LLIAM and TimeLLM is the average RMSE obtained after aggregating the test instances.
@@
of Avgs.
LLIAM T=10 0.0981 0.1615 0.1706 0.1472 0.1023 0.2400 0.1146 0.1661 0.0175 0.1353
LLIAM T=20 0.0944 0.1616 0.1743 0.1495 0.1024 0.2486 0.1178 0.1609 0.0168 0.1363
TimeLLM 0.1339 0.3483 0.4941 0.1458 0.0964 0.2350 0.1155 0.1281 0.0200 0.1908
GRU 0.2410 0.6040 0.6160 0.1650 0.1330 0.2590 0.1350 0.4310 0.0300 0.2904
LSTM 0.2390 0.6240 0.5790 0.1530 0.1020 0.2470 0.1270 0.4270 0.0320 0.2811
TCN D=0 0.2420 0.4360 0.4810 0.1600 0.1020 0.2460 0.1290 0.2100 0.0350 0.2268
TCN D=0.1 0.2320 0.4710 0.4850 0.1660 0.1110 0.2490 0.1290 0.2300 0.0340 0.2341
TCN D=0.25 0.2470 0.4780 0.5330 0.1680 0.1080 0.2570 0.1320 - - 0.2747
Table 6: SMAPE metric per model and dataset. LSTM, GRU and TCN SMAPE are averages for all the configurations tested, meanwhile LLIAM and TimeLLM is the average SMAPE obtained after aggregating the test instances.
@@
ILI Weather Average
LLIAM T=10 0.000 0.000 0.493 0.000 0.000 14.414 0.000 0.000 1.313 1.802
LLIAM T=20 0.000 0.000 0.493 0.000 0.000 7.207 0.000 0.000 0.263 0.885
Table 7: Percentage of test instances unable to decode into the expected forecast horizon. GRU, LSTM, TCN and TimeLLM models have been omitted because their missing rate is always 0%.
Figure 3: Critical differences diagram with Wilcoxon post-hoc analysis using the SMAPE metric at $\alpha=0.05$. Models connected by a horizontal line indicate no statistically significant difference detected by the Wilcoxon test between them. A rank for each model is also indicated next to it.
One relevant aspect to highlight is the ability of our proposal to produce coherent and expected responses. It must be able to predict at any given time without resorting to making up anomalies or the so-called LLM hallucinations. The missing rate for each dataset and LLIAM configuration is reported in Table 7. Hallucinations were observed on three of the nine datasets: NN5 Daily with the highest rate (14.41% with T=10 and 7.207% with T=20) followed by Weather (1.313% with T=10 and 0.263% with T=20) and lastly M1 Quarterly (0.493% for both T=10 and T=20). The most frequent anomaly is due to the model stopping generating tokens early, failing to reach the expected forecast horizon. Early studies showed higher missing rates, but increasing the forecast horizon by one step ($h+1$) reduced this issue. Overall, the total hallucinations are not significant with respect to the results obtained, achieving an average lower than 2% with T=10 and lower than 1% with T=20. As method based on LLM, TimeLLM also generates hallucinations, as is common in such models, but it cannot detect and deal with them properly as they are absorbed by the final projection layer present in its architecture.
In regard to the computational cost related to the training of these models, experiments of the methods based on RNNs and TCNs required extensive training times, as all combinations of hyperparameters have to be tried for every dataset. TimeLLM also employs a dataset-by-dataset training procedure, but requires much less training time because it does not require hyperparameter optimization. LLIAM benefits more from Transfer Learning, enabling the training of a model across all datasets on a single GPU in approximately the same time as TimeLLM. This emphasizes the efficiency of our method in reducing training overhead while maintaining generalization across datasets.
4.3 Zero-shot study
In this second study how LLIAM performs on datasets with unknown domains and distributions is tested. We also want to check how the integration of TSF expertise affects the model by comparing the results obtained with respect to its base model, LLaMA-1 7B. TimeLLM could not be included in this study because its architecture has to be trained independently for each dataset, due to an output projection that depends on the number of instants to predict as explained in Section 4.2.
Methodology
The weights of LLIAM after training with the comparative study datasets and the LLAMA-1 7B weights provided by Meta are used for inference only. The test partitions of the ETTh1, ETTh2 and San Francisco Traffic datasets are employed, constructed in the same way as explained in the experimental study and applying the anomalies mitigation strategy described above (Section 4.2). Due to the large length of the ETTh1 and ETTh2 datasets, the last 10% of the generated windows to test the models are considered. Each model is evaluated using T=10 and T=20 and the prompting scheme proposed in Section 3 is employed in both base model and LLIAM to evaluate only the impact of the new knowledge incorporated over LLIAM with respect LLaMA-1 7B.
Results
Table 8 highlights the average SMAPE and RMSE obtained for each model configuration evaluated. Attention may be drawn by the high average error obtained with LLaMA-1 7B T=20 and T=10 in the dataset ETTh1. This is produced by an anomalous behavior, where the model predicted a very high value in some test windows lags whose tendency is led by low ones, explaining why the SMAPE tends to be much lower. We assume that this happens because LLaMA, although it is capable of emitting forecasts, it does not have the specific knowledge integrated in LLIAM and it may not achieve the intended objective of the task. LLIAM is more robust in comparison with its base model. Generalization on unknown domains is slightly better when using LLIAM instead of LLaMA can also be affirmed, supported by the results obtained. The use of a common prompt template during the training phase of LLIAM achieves its objective of letting the model know when it is facing TSF tasks.
RMSE (Lower is better) SMAPE (Lower is better)
ETTh1 ETTh2
San Francisco
Traffic
ETTh1 ETTh2
San Francisco
Traffic
@@
LLIAM T=10 2.048 10.176 1.598 0.203 0.236 0.131 0.190
LLIAM T=20 2.052 9.837 1.648 0.216 0.228 0.137 0.194
LLAMA T=10 1926800.126 701.422 158.113 0.252 0.236 0.368 0.285
LLAMA T=20 1.93e+111 667.916 159.401 0.247 0.234 0.372 0.284
@@
5 Conclusions
This article introduces a straightforward and general approach for adapting FMs, in particular LLMs, to the TSF task. LLIAM combines the PEFT technique LoRA with a simplification of the PromptCast prompting scheme to improve the performance of LLaMA by integrating task-specific knowledge into the model without increasing its number of parameters. The textual data representation of the input that characterizes LLMs is respected in the application of the prompting scheme. This also serves as a differentiator, helping the LLM to identify when it faces a TSF task. A two-stage experimentation was conducted to demonstrate the capabilities of LLIAM over RNNs, TCNs, another state-of-the-art LLM-base approach, TimeLLM, and the base model employed, LLaMA.
In the comparative study it was demonstrated that LLM-based approaches achieve a better performance than the conventional DL ones without the need for intense hyperparameter-tuning. This find suggests that there is a certain similarity between language and time-series that LLM-based models are able to exploit, taking advantage of the knowledge transfer that comes from being pre-trained with a large and heterogeneous amount of data. The straightforward modifications applied to LLIAM are sufficient for enhancing the capabilities of LLMs, as it achieves similar behaviour to TimeLLM and outperforms it with a fairer metric such as SMAPE. Furthermore, the dataset-specific training of TimeLLM, its reliance on contextual information, and its pre-processing are a disadvantage with respect to the one-pass approach of LLIAM. This demonstration serves to illustrate that the methodology employed in this study produces a more general model. The results obtained in the zero-shot study confirm to us that the addition of specialized knowledge related to TSF problems enhances the performance of LLaMA-1 7B, highlighting its generalization capacity over unknown datasets and the robustness of the predictions emitted. TimeLLM is not compatible with the zero-shot study design used in this paper because it needs to be trained with information from the dataset to be predicted in addition to the limitations of its architecture.
The findings of the present study have provided answers to the questions that were raised at the initiation of the research. It has been demonstrated that LLMs are capable of leveraging the similarity of text and time-series by implementing easy modifications to them, achieving more general and straightforward solutions than other methods.
This work sets the stage for future research involving more extensive testing with different LLMs, including adaptations or architectures specifically tailored for TSF tasks and using larger models, but respecting the original architecture of these models and trying to modify it as little as possible. In addition, the exploration and development of PEFT techniques beyond LoRAs is a promising direction for improving the efficiency and adaptability of these models. Another field to explore is how time-series data have to be inputted into this type of models, including determining what dataset and domain specific information is needed to activate the capabilities of these models as time-series forecasters. Overall, the arrival of the FMs opens up the possibility of numerous interesting proposals.